40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4079     Entry time: Mon Dec 20 23:10:25 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Pretty much ready for pump-down. A few final things.... 

[Kiwamu, Jenne, Koji, Osamu]

We have mostly prepared the IFO for pump down. 

After lunch [Steve, Bob, Koji, Kiwamu, Jenne, Joe, Joon Ho, Vladimir, Osamu] put the access connector back in place.  Hooray!  Steve still has to check the Jam Nuts before we pump down.  Kiwamu checked the leveling of the IOO table, and fixed all of the weights to the table.

For all 4 test masses, bars (upside-down dog clamps) were placed to mark the alignment of 2 sides of the suspension tower.  All test mass tables were re-leveled, and the weights fixed to the tables.  

For ETMY, PRM, BS, SRM, we confirmed that the OSEMs were close to their half-range.  ETMX was already fine.  ITMY (the screens and the optics wiki are still old-convention, so this is listed as ITMX! No good!) OSEMs are pretty much fine, but ITMX desperately needs to be adjusted.  Unfortunately, no one can find the standard screwdriver (looks like a minus), to adjust the ITM OSEMs.  All the other towers had hex-key set screws, but the ITMs need a screwdriver.  We will ask Bob to sonicate a screwdriver in the morning. 

 

 

ELOG V3.1.3-